# INDICT: Code Generation with Internal Dialogues of Critiques for Both Security and Helpfulness

### v1

@@ Appendix E Details of Experimental Results
Appendix F Instruction Prompts
We described the example instruction prompts we used in this section. For each prompt template, depending on the model roles and tasks, we replace the following placeholders with applicable input components: {question} and {answer} are replaced with the corresponding task description and latest model output from the actor LLM. During the posthoc feedback stage, {answer} is also concatenated with any execution results (e.g. test outcomes, error messages) after executing the corresponding extracted code output with a code interpreter. {scratchpad} is typically used as a placeholder to contain past interactions between the two critics.
@@ Appendix F Instruction Prompts
Listing 4: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeSearch actions.
@@
Listing 5: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeReview actions.
@@
Listing 6: Prompt with instruction for a critic model to focus on safety-related concerns in open-ended generation tasks.
@@
Listing 7: Prompt with instruction for a critic model to focus on helpfulness-related concerns in open-ended generation tasks.
@@
Listing 8: Prompt with instruction for a summarizer model to summarize the interactions between safety and helpfulness critics for open-ended generation tasks.

### v2

@@ Appendix H Details of Experimental Results
Appendix I Instruction Prompts
We described the example instruction prompts we used in this section (Listing 1 to 10). For each prompt template, depending on the model roles and tasks, we replace the following placeholders with applicable input components: {question} and {answer} are replaced with the corresponding task description and latest model output from the actor LLM. During the posthoc feedback stage, {answer} is also concatenated with any execution results (e.g. test outcomes, error messages) after executing the corresponding extracted code output with a code interpreter. {scratchpad} is typically used as a placeholder to contain past interactions between the two critics.
Note that INDICT uses zero-shot prompting in each step. We prompt the critic agent to condition the current critique and generate a unique query to obtain more knowledge. We extract the search keywords following our instruction templates e.g. in the form of ’Search[keyword]’. For generating code snippets, we prompt the model similarly but ask the model to wrap the output code in “‘. Note that the current prompts in the Appendix are for tool-finetuned models like CommandR which automatically generates tool parameters given a tool calling function definition.
@@ Appendix I Instruction Prompts
Listing 4: Prompt with instruction for critic models to generate relevant parameters (text queries) to perform CodeSearch/CodeReview actions.
You are an advanced intelligent agent with direct access to Internet. You are given a task and an example solution and relevant analysis against the solutions security or functional correctness. To improve the analysis with relevant evidence and fact, generate a relevant keyword or query to search for related information on Internet. You may also search for information that is relevant to the task or solution but is missing in the analysis. Use the following format: Search[<query or keyword>].
Task: {question}
Solution: {answer}
{scratchpad}
Query (in the form of Search[<query or keyword>]):
Listing 5: Prompt with instruction for critic models to generate relevant parameters (code snippets) to perform CodeSearch/CodeReview actions.
You are an advanced intelligent agent with direct access to Internet. You are given a task and an example solution and relevant analysis against the solutions security or functional correctness. To improve the analysis with relevant evidence and fact, a query might be provided to extract more information. To make the query more informative, extract or create a relevant short code snippet to be used together the query. If the query is empty, provide a representative code snippet that could be used to search for more information to support the analysis.
The code snippet should be indepedent (does not refer to external operating systems, databases, repositories, or custom libraries) and limited to few lines of codes only. Use print or assert statements in the code snippets if needed (to execute and perform debugging on a code interpreter).
Wrap the code snippet in ‘‘‘.
Task: {question}
Solution: {answer}
{scratchpad}
Query: {query}
Short code snippet in a single code block (wrap in ‘‘‘):
Listing 6: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeSearch actions. This prompt is specifically for CommandR models which by default can receive any predefined external tools and select relevant ones given the context of the prompt.
@@
Listing 7: Prompt with instruction for critic models to generate relevant parameters (text queries and/or code snippets) to perform CodeReview actions. This prompt is specifically for CommandR models which by default can receive any predefined external tools and select relevant ones given the context of the prompt.
@@
Listing 8: Prompt with instruction for a critic model to focus on safety-related concerns in open-ended generation tasks.
@@
Listing 9: Prompt with instruction for a critic model to focus on helpfulness-related concerns in open-ended generation tasks.
@@
Listing 10: Prompt with instruction for a summarizer model to summarize the interactions between safety and helpfulness critics for open-ended generation tasks.
@@
Appendix J Qualitative Analysis
To explain the benefits of INDICT and where INDICT may fail, we included example generation outputs from INDICT and related methods. In Figure 9, we show that given a code generation task, INDICT can generate code that is more secure and robust than strong baselines (Direct generation, Reflexion [Shinn et al., 2023], and CAMEL [Li et al., 2024]). In Figure 10, we illustrate cases where INDICT may fail due to nontrivial errors.
Compared to the baseline methods, our approach can simultaneously improve both the helpfulness and safety of the output code generation. Specifically, given a relevant information snippet by the safety critic (about the hashing method SHA-256), our actor agent correctly revised the code with a more secure hashing method, avoiding using MD5 hashing and the common security risk CWE-328. At the same time, our generated code is generally more helpful with properly modularized functions implementing supporting features such as input validations. As we noted, this feature has generally emerged in code solutions by collaborative agent systems like CAMEL and INDICT.
Figure 9: Qualitative comparison between our method and related baselines
Figure 10: Qualitative analysis of failure cases when applying our method